THE FOURIER TRANSFORM MICROWAVE/MILLIMETER SPECTRUM OF ScO (X^{2}Σ^{+})

\underline{D. T. HALFEN}, J. MIN, and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectra of ScO (X^{2}Σ^{+}) have been recorded in the 4 - 90 GHz range using Fourier transform microwave/millimeter (FTM/mmW) techniques. This species was created in a supersonic jet expansion of laser-ablated scandium vapor and N_{2}O gas, diluted in argon. The N = 1 → 0 and 2 → 1 rotational transitions in both v = 0 and 1 have been measured near 30 and 61 GHz, respectively. The data over 60 GHz were obtained using a new E-band (60 - 90 GHz) FTmmW spectrometer system. The data have been analyzed, and rotational, fine, and hyperfine constants have been determined, which are in good agreement with those from past optical studies. ScO is a potential circumstellar molecule in giant/supergiant stars, where it is produced in oxygen-burning nucleosynthesis.